### v1

@@ Abstract
Video generation techniques have made remarkable progress, promising to be the foundation of interactive world exploration. However, existing video generation datasets are not well-suited for world exploration training as they suffer from some limitations: limited locations, short duration, static scenes, and a lack of annotations about exploration and the world. In this paper, we introduce Sekai (meaning “world” in Japanese), a high-quality first-person view worldwide video dataset with rich annotations for world exploration. It consists of over 5,000 hours of walking or drone view (FPV and UVA) videos from over 100 countries and regions across 750 cities. We develop an efficient and effective toolbox to collect, pre-process and annotate videos with location, scene, weather, crowd density, captions, and camera trajectories. Experiments demonstrate the quality of the dataset. And, we use a subset to train an interactive video world exploration model, named YUME (meaning “dream” in Japanese). We believe Sekai will benefit the area of video generation and world exploration, and motivate valuable applications.
We are looking for collaboration and self-motivated interns interested in interactive world generation. Contact: zhangkaipeng@pjlab.org.cn
@@ 1 Introduction
Recent advancements in video generation zhang2025packing ; yang2024cogvideox ; opensora2 ; wan2025wan ; kong2024hunyuanvideo have been remarkable, making it a promising approach for world generation through video generation. Meanwhile, camera-controlled video generation hecameractrl ; namekata2024sg ; hou2024training is a suitable way for world exploration, since camera trajectories can be converted by keyboard and mouse inputs. However, generating long and realistic videos with precise camera control remains a significant challenge. Data is a critical challenge. Existing video generation datasets nan2024openvid ; chen2024panda ; ju2024miradata are not well-suited for world exploration as they suffer from limitations: limited locations, short duration, static scenes, and a lack of annotations about exploration (e.g., camera trajectories) and world annotations (e.g., location, weather and scene).
[…]
We do dataset statistics analysis to show the scale and diversity of the dataset. And, we validate the accuracy of annotations for YouTube videos. Besides, we use a subset of the Sekai-Real to train an interactive video world exploration model, named yume ({CJK}UTF8minゆめ, meaning “dream” in Japanese).
[…]
The main contribution of this work is the dataset, Sekai. It is a high-quality long-form video dataset for worldwide exploration by walking or drone. We annotate diverse labels, including captions, location, scene, weather, crowd density, time-of-day, and camera trajectories for each video. Experiments demonstrate the effectiveness and the quality of the dataset. We believe Sekai will benefit the areas of video generation and world generation, and motivate valuable applications. We provide an introduction video on the project page.
@@ 2.1 World Generation Model
Recent years have seen a growing interest in video generation zhang2025packing ; kong2024hunyuanvideo ; wang2025wan ; chen2024videocrafter2 ; li2024t2v ; xing2024dynamicrafter ; xu2024easyanimate , 3D scene generation yu2024wonderworld ; yu2024wonderjourney ; chung2023luciddreamer ; engstler2024invisible ; hollein2023text2room ; zhang2024text2nerf , and 4D generation bahmani20244d ; bahmani2024tc4d ; zhang20244diffusion ; xu2024comp4d ; zhang2025matrixgame , with significant advancements opening up new possibilities in the development of world generation models brooks2024video ; agarwal2025cosmos ; kang2024far ; meng2024towards ; xiang2024pandora . In the realm of video generation, text-to-video generation chen2024videocrafter2 ; li2024t2v has played a pivotal role, achieving high-fidelity results, while image-to-video generation xing2024dynamicrafter ; xu2024easyanimate ; shi2024motion has also seen notable advancements. Sora brooks2024video further underscores the significance of video generation in the context of world generation models. Among 3D scene generation methods, techniques chung2023luciddreamer ; engstler2024invisible ; hollein2023text2room ; yu2024wonderworld utilize depth estimation models li2024megasam ; chen2025video ; piccinelli2024unidepth to extend 2D scenes into 3D representations. 4D scene generation bahmani20244d ; bahmani2024tc4d ; zhang20244diffusion further introduces dynamics, focusing on the evolution of objects or scenes over time xu2024comp4d and dynamic interactions zhang2025matrixgame . This paper primarily focuses on interactive video generation for world exploration, aiming to construct a dynamic and realistic world using image, text, or video for unrestricted exploration.
@@ 3 Dataset Curation
Figure 3: The dataset curation pipeline. * indicates we annotate trajectories for partial data.
@@ 3.2 Video Pre-processing
Camera Trajectory Filtering. For Sekai-Real, we employ a state-of-the-art structure from motion (SfM) model to extract camera trajectories. However, some trajectories exhibit implausible or counter-intuitive motions, so we heuristically filter out abnormal cases using the following rules. Specifically, we exclude video clips if they satisfy either of the following: (1) Multiple abrupt trajectory reversals (i.e., directional changes exceeding 150 degrees) within a 10-second window. (2) A camera viewpoint shift greater than 60 degrees between two consecutive frames. (3) A camera position displacement greater than 5 times the average displacement of the 30 consecutive frames containing these two frames. This filtering phase is only conducted for partial data annotated with camera trajectories.
@@ 3.3 Video Annotation of Sekai-Real
We annotate video clips using large vision-language models.
@@
Camera Trajectories. We experimented with various camera trajectory annotation methods of different types, including the visual odometry method DPVO teed2023deep , the deep visual SLAM framework MegaSaM li2024megasam , and a carefully designed 3D transformer VGGT wang2025vggt that outputs 3D quantities. Through empirical experiments and comparisons, we chose MegaSaM as the baseline annotation method and made adjustments to optimize annotation accuracy and efficiency. First, we replaced the monocular depth estimation model Depth Anything yang2024depth used in MegaSaM with Video Depth Anything chen2025video , which performs better in terms of temporal consistency. We also tried modifying the metric depth estimation model in MegaSaM by replacing UniDepth piccinelli2024unidepth with DepthPro bochkovskii2024depth . Additionally, we optimized the official implementation of MegaSaM to support cross-machine, multi-GPU parallel inference, significantly improving annotation efficiency. We annotate over 600 hours of videos (after the sampling).
@@ 3.5 Video Sampling
Given the prohibitive cost of training on the full Sekai-Real, we propose a strategy to sample the best of the best clips with the highest quality and diversity. The number is related to the computational budget for further video generation model training. In this paper, we sample 400 hours of the best of the best videos as Sekai-Real-HQ.
@@ 3.5.2 Diversity Sampling
We balance the videos using the following modules one by one. And for Sekai-Real-HQ, the sampling ratio $\alpha_{content}$, $\alpha_{loc}$, $\alpha_{cate}$, $\alpha_{camera}$ are equal to $70\%$, $60\%$, $60\%$, and $75\%$, respectivelty.
@@ 4 Dataset Statistics
For the Sekai-Game collection, we have considered the data balance issue while playing the game. Figure 4 provides an overview of the statistical information of the Sekai-Real dataset. Overall, the Sekai-Real dataset includes 101 countries and regions, with the video durations exhibiting a distinct long-tail distribution. The top eight countries (such as the Japan, the United States, and the United Kingdom) account for approximately 60% of the total duration. From various perspectives, we categorize the Sekai-Real dataset into four weather types, four scene types, four time-of-day categories, and five levels of crowd density. Specifically, the majority of the videos were outdoors. For outdoor videos, the weather conditions are predominantly sunny and cloudy, and the dataset also includes unique weather situations, such as rain and snow, increasing the diversity of the dataset. Regarding the time of filming, most videos were shot during the daytime, followed by nighttime footage. This provides a mix of well-lit scenes under natural light and darker scenes with artificial lighting, offering diverse challenges for model learning. In terms of crowd density, the distribution is relatively uniform, ranging from the sparsely populated forests of Finland to the bustling streets of Japan. This stepwise distribution covers a variety of crowd density scenarios, providing diverse training settings for the model, such as curriculum learning and evaluation of video generation models under varying levels of crowd density.
@@
The statistics of Sekai-Real and Sekai-Real-HQ across multiple dimensions are shown in Figure 5. Sekai-Real-HQ is the best of the best subset of Sekai-Real, and with more balanced data distribution. Seeing Figure 5 (a), Sekai-Real demonstrates strong overall video quality scores, with more than half of the videos scoring above 0.9, while Sekai-Real-HQ exhibits a higher mean video quality score and a lower variance to address the long-tail distribution issue. Figure 5 (b) shows the distribution of Sekai-Real’s locations. Both Sekai-Real and Sekai-Real-HQ cover a wide range of countries globally. Sekai-Real-HQ demonstrates a more balanced distribution, which is more effective in mitigating potential bias during model training. Figure 5 (c) shows the distribution of caption token counts in Sekai-Real and OpenVid-1M. It can be observed that the average number of caption tokens in Sekai-Real exceeds 200, which is significantly higher than OpenVid-1M’s average of around 130. Figure 5 (d) shows the distribution of camera trajectory jitter values before and after applying Camera Trajectory Diversity. We can observe that the distribution for Sekai-Real-HQ is smoother compared to that of the Subset of Sekai-Real (500 hours, before applying Camera Trajectory Diversity). This indicates that Sekai-Real-HQ achieves better diversity and a more uniform distribution.
5 YUME Model
Figure 6: Examples of videos generated by YUME using keyboard and mouse control.
We train an interactive world exploration model named YUME ({CJK}UTF8minゆめ, meaning “dream” in Japanese) using a subset of the Sekai-Real-HQ. Specifically, it receives an image and allows unrestricted exploitation using keyboard and mouse control from users. We show some examples in Figure 6.

### v3

@@ Abstract
Video generation techniques have made remarkable progress, promising to be the foundation of interactive world exploration. However, existing video generation datasets are not well-suited for world exploration training as they suffer from some limitations: limited locations, short duration, static scenes, and a lack of annotations about exploration and the world. In this paper, we introduce Sekai (meaning “world” in Japanese), a high-quality first-person view worldwide video dataset with rich annotations for world exploration. It consists of over 5,000 hours of walking or drone view (FPV and UVA) videos from over 100 countries and regions across 750 cities. We develop an efficient and effective toolbox to collect, pre-process and annotate videos with location, scene, weather, crowd density, captions, and camera trajectories. Comprehensive analyses and experiments demonstrate the dataset’s scale, diversity, annotation quality, and effectiveness for training video generation models. We believe Sekai will benefit the area of video generation and world exploration, and motivate valuable applications.
@@ 1 Introduction
Recent advancements in video generation zhang2025packing; yang2024cogvideox; opensora2; wan2025wan; kong2024hunyuanvideo have been remarkable, making it a promising approach for world generation through video generation. Meanwhile, camera-controlled video generation hecameractrl; namekata2024sg; hou2024training is a suitable way for world exploration, since camera trajectories can be converted by keyboard and mouse inputs. However, generating long and realistic videos with precise camera control remains a significant challenge. A major bottleneck lies in the data itself. Existing video generation datasets nan2024openvid; chen2024panda; ju2024miradata are not well-suited for world exploration as they suffer from limitations: limited locations, short duration, static scenes, and a lack of annotations about exploration (e.g., camera trajectories) and world annotations (e.g., location, weather and scene).
In this paper, we introduce Sekai (せかい, meaning “world” in Japanese), a high-quality egocentric worldwide video dataset for world exploration (see Figure 1 and Figure 2). Most videos contain audio for an immersive world generation. It also benefits other applications, such as video understanding, navigation, and video-audio co-generation. Sekai-Real comprises over 5000 hours of videos collected from YouTube with high-quality annotations. Sekai-Game comprises videos from a realistic video game, Lushfoil Photography Sim, with ground-truth annotations. It has five distinct features: (1) High-quality and diverse video. All videos are recorded in 720p at 30 FPS, featuring diverse weather conditions, various times, and dynamic scenes. (2) Worldwide location. Videos are captured across 101 countries and regions, featuring over 750 cities with diverse cultures, activities, architectures, and landscapes. (3) Walking and drone view. Beyond the walking videos (e.g., citywalk and hiking), Seikai contains drone view (FPV and UAV) videos for unrestricted world exploration. (4) Long duration. All walking videos are at least 60 seconds long, ensuring real-world, long-term world exploration. (5) Rich annotations. All videos are annotated with location, scene, weather, crowd density, captions, and camera trajectories. YouTube videos’ annotations are of high quality, while annotations from the game are considered ground truth.
To construct the Sekai dataset, we develop a curation pipeline (see Section 3) for Sekai-Real (YouTube videos) and Sekai-Game (video game videos). (1) For Sekai-Real, we first manually search and download high-quality walking and drone videos. Then we introduce a pre-processing pipeline to obtain video clips by shot detection, video transcoding, and quality evaluation. After that, we develop an annotation framework to annotate location, scene type, weather, crowd density, captions, and camera trajectories. Considering the large amount of data and practical usage, we further introduce a video sampling module to sample the top-tier videos according to the computational resources for training the video generation model. (2) For the Sekai-Game, we first play Lushfoil Photography Sim and record videos. Then we use the same pre-processing pipeline to obtain video clips. For the annotation, we develop a toolbox to record ground-truth annotations while playing.
We conduct statistical analyses to characterize the scale and diversity of the dataset and independently validate the accuracy of YouTube annotations. We then fine-tune a video generation foundation model on the the top tier of Sekai-Real for text-to-video and image-to-video, yielding consistent gains in world-exploration scenarios, especially in video dynamics and visual quality. In addition, leveraging Sekai’s camera trajectory annotations, we train for interactive video generation, where the model takes a camera trajectory as input and generates videos consistent with the intended camera motion. Across Sekai-Real and Sekai-Game, this training substantially improves interaction following, significantly reducing the error between the trajectories of the generated video and the target.
Figure 1: Sekai is collected from Youtube and a video game. It consists of walking and drone-view egocentric videos with recorded audio. We provide rich annotations of camera trajectories, location, crowd density, scene, weather, time of day, and captions.
To summarize, our contributions are threefold:
  • We introduce Sekai, a large-scale, high-quality long-form video dataset for worldwide exploration via walking and drone footage, with rich annotations.
  • We develop a curation pipeline that efficiently collects, filters, and annotates videos from the web and from video games.
  • We validate the quality and effectiveness of the dataset through comprehensive analyses, annotation verification, and experiments on various video generation tasks.
@@ 2.1 World Generation Model
Recent years have seen a growing interest in video generation zhang2025packing; kong2024hunyuanvideo; wan2025wan; chen2024videocrafter2; li2024t2v; xing2024dynamicrafter; xu2024easyanimate, 3D scene generation yu2024wonderworld; yu2024wonderjourney; chung2023luciddreamer; engstler2024invisible; hollein2023text2room; zhang2024text2nerf, and 4D generation bahmani20244d; bahmani2024tc4d; zhang20244diffusion; xu2024comp4d; zhang2025matrixgame, with significant advancements opening up new possibilities in the development of world generation models brooks2024video; agarwal2025cosmos; kang2024far; meng2024towards; xiang2024pandora. In the realm of video generation, text-to-video generation chen2024videocrafter2; li2024t2v has played a pivotal role, achieving high-fidelity results, while image-to-video generation xing2024dynamicrafter; xu2024easyanimate; shi2024motion has also seen notable advancements. Sora brooks2024video further underscores the significance of video generation in the context of world generation models. Among 3D scene generation methods, techniques chung2023luciddreamer; engstler2024invisible; hollein2023text2room; yu2024wonderworld utilize depth estimation models li2024megasam; chen2025video; piccinelli2024unidepth to extend 2D scenes into 3D representations. 4D scene generation bahmani20244d; bahmani2024tc4d; zhang20244diffusion further introduces dynamics, focusing on the evolution of objects or scenes over time xu2024comp4d and dynamic interactions zhang2025matrixgame. This paper primarily focuses on interactive video generation for world exploration, aiming to construct a dynamic and realistic world using image, text, or video for unrestricted exploration.
@@ 3 Dataset Curation
Figure 3: The dataset curation pipeline. *indicates that the statistics were derived from a subset of trajectory annotations.
@@ 3.2 Video Pre-processing
Camera Trajectory Filtering. For Sekai-Real, we employ a state-of-the-art structure from motion (SfM) model to extract camera trajectories. However, some trajectories exhibit implausible or counter-intuitive motions, so we heuristically filter out abnormal cases using the following rules. Specifically, we exclude video clips if they satisfy either of the following: (1) Multiple abrupt trajectory reversals (i.e., directional changes exceeding 150 degrees) within a 10-second window. (2) A camera viewpoint shift greater than 60 degrees between two consecutive frames. (3) A camera position displacement greater than 5 times the average displacement of the 30 consecutive frames containing these two frames. This filtering phase is performed on partially selected data annotated with trajectories.
@@ 3.3 Video Annotation of Sekai-Real
We annotate the video data from multiple perspectives, including geographic locations, content category and caption, and per-frame camera trajectories.
@@
Camera Trajectories. We experiment with various camera trajectory annotation methods of different types, including the visual odometry method DPVO teed2023deep, the deep visual SLAM framework MegaSaM li2024megasam, and a carefully designed 3D transformer VGGT wang2025vggt that outputs 3D quantities. Through empirical experiments and comparisons, we choose MegaSaM as the baseline annotation method and made adjustments to optimize annotation accuracy and efficiency. Additionally, we replace the monocular depth estimation model Depth Anything yang2024depth used in MegaSaM with Video Depth Anything chen2025video, which performs better in terms of temporal consistency. We also optimize the official implementation of MegaSaM to support cross-machine, multi-GPU parallel inference, significantly improving annotation efficiency.
@@ 3.5 Video Sampling
Given the prohibitive cost of training on the full Sekai-Real, we propose a strategy to sample the top-tier clips with the highest quality and diversity. The number is related to the computational budget for further video generation model training. In this paper, we sample 400 hours of the videos as Sekai-Real-HQ.
@@ 3.5.2 Diversity Sampling
We balance the videos using the following modules one by one. And for Sekai-Real-HQ, the sampling ratio $\alpha_{content}$, $\alpha_{loc}$, $\alpha_{cate}$, $\alpha_{camera}$ are equal to $70\%$, $\%, 6060\%$, and $75\%$, respectivelty.
@@ 4 Dataset Statistics
Figure 4 summarizes the statistics of Sekai-Real, which covers 101 countries and regions with a clear long-tail distribution in video duration. The top eight countries (e.g., Japan, the United States, and the United Kingdom) account for about 60% of the total duration. The dataset is categorized by four weather types, four scene types, four time-of-day categories, and five crowd-density levels from various perspectives. Specifically, most videos are outdoor scenes, primarily under sunny or cloudy conditions, while rain and snow further enrich diversity. Daytime footage dominates, followed by nighttime scenes, providing a range of lighting conditions for model learning. Crowd density is evenly distributed, from sparse rural areas to densely populated city streets, supporting tasks such as curriculum learning and evaluation under varying crowd levels. For the Sekai-Game collection, data balance was considered during gameplay.
@@
The statistics of Sekai-Real and Sekai-Real-HQ across multiple dimensions are shown in Figure 5. Sekai-Real-HQ, a top-tier subset of Sekai-Real, features a more balanced data distribution. Seeing Figure 5 (a), Sekai-Real demonstrates strong overall video quality scores, with more than half of the videos scoring above 0.9, while Sekai-Real-HQ exhibits a higher mean video quality score and a lower variance to address the long-tail distribution issue. Figure 5 (b) shows the distribution of Sekai-Real’s locations. Both Sekai-Real and Sekai-Real-HQ cover a wide range of countries globally. Sekai-Real-HQ demonstrates a more balanced distribution, which is more effective in mitigating potential bias during model training. In terms of captions, Figure 5(c) shows that Sekai-Real exhibits a higher average token count compared to OpenVid-1M nan2024openvid, providing richer textual supervision. Figure 5 (d) shows the distribution of camera trajectory jitter before and after applying camera trajectory sampling. We can observe that the distribution for Sekai-Real-HQ is smoother than that of partially selected Sekai-Real data. This indicates that Sekai-Real-HQ achieves better diversity and a more uniform distribution.
5 Experiments
In this section, we first validate the quality of the annotation in Sekai. Then we use the top-tier subset Sekai-Real-HQ to fine-tune a video generation foundation model for text-to-video generation and image-to-video generation to validate the effectiveness of the data. Additionally, we explore interactive video generation with the camera trajectories annotated in Sekai.
5.1 Evaluation of Annotation Quality
We evaluate the quality of the annotated location, camera trajectory, and category.
Location Quality. For videos in Sekai-Real, formatted locations were standardized using GPT-4o hurst2024gpt based on original YouTube titles and descriptions. To evaluate their quality, we randomly sampled 500 videos and asked co-authors to verify each location using online maps, checking for three possible issues: (1) Omission – missing or incorrectly merged segments; (2) Temporal mismatch – misalignment between the timestamp and location; (3) Location hallucination – inferred rather than explicitly stated locations. Since the title and description provide rich information for GPT-4o, the evaluation revealed that such issues were rare (<5%), indicating that the overall location quality produced by GPT-4o was remarkably high.
Camera Trajectory Quality. We use two representative methods of two categories in the field of Structure from Motion, cascaded and end-to-end, namely, MegaSAM li2024megasam and VGGT wang2025vggt, to annotate 100 walking video clips and 100 drone-view video clips to evaluate their accuracy of camera trajectory prediction. Our validation experiments show that: (1) MegaSAM produces smoother camera trajectories than VGGT, as it has a global optimization module. (2) VGGT offers the advantage of faster inference speed but lower annotation quality. Since data quality is our priority and smoother camera trajectories are more beneficial for model training, we opted for MegaSAM to annotate camera trajectories.
Category Quality. We randomly sample 500 examples and ask co-authors to label them for evaluating category annotation quality. The sampling strategy follows the category diversity sampling to ensure label variety. The results show that the overall agreement between Qwen2.5-VL and human annotations exceeds 90%. For weather, most discrepancies occur between cloudy/foggy and rainy labels. The difficulty lies in that rain is often imperceptible in a single frame and becomes evident only across consecutive frames, indicating that large vision-language models still struggle with temporal visual reasoning. We plan to leverage dedicated weather prediction datasets to fine-tune video understanding models wang2024internvideo2 for more accurate weather annotations. Notably, we exclude indoor scenes from the analysis, as indoor lighting often affects the accuracy of both weather and time-of-day predictions for models and human annotators.
5.2 Video Generation
Table 1: Evaluation across training steps for text-to-video generation and image-to-video generation. Higher is better for all metrics.
[…]
We fine-tune the SkyReels-V2 chen2025skyreels video generation foundation model on the Sekai-Real-HQ dataset for text-to-video generation and image-to-video generation to validate the effectiveness of the dataset.
5.2.1 Settings
[…]
Evaluation Dataset. For a fair evaluation, Sekai-Real-HQ is randomly divided into ten folds, with the last two folds used as the candidate test set and excluded from the training set. Subsequently, independent annotators are invited to manually select 50 clips from the candidate test set, with an emphasis on maintaining diversity during the selection process.
Evaluation Metrics. We adopt the VBench huang2024vbench evaluation metrics to comprehensively assess the model’s performance at different training steps. Specifically, Subject Consistency measures whether the subject’s appearance remains consistent. Background Consistency evaluates the temporal stability of background scenes. Motion Smoothness assesses whether motion is smooth and physically plausible. Dynamic Degree quantifies the extent of motion to avoid static videos. Aesthetic Quality reflects the perceived artistic and visual appeal. Finally, Imaging Quality evaluates distortions such as over-exposure, noise, and blur. For the image-to-video generation task, we further adopt the VBench++ huang2024vbench++ metrics I2V Subject and I2V Background to better evaluate the alignment between the prompt image and the generated video.
5.2.2 Quantitative Results
Table 1 presents the results of fine-tuning video generation foundation models in different tasks on the Sekai-Real-HQ. We observed that (1) on both text-to-image generation and image-to-video generation tasks, the overall generation quality improves with training, with consistent gains across most metrics and a steady rise in the Overall Score. (2) For image-to-video generation, the I2V Subject and I2V Background metrics improve markedly, imaging quality increases steadily through training, and the Overall Score keeps rising. (3) Dynamic Degree decreases after early peaks. This is a rebalancing where the model shifts from initially exaggerating motion to focusing more on sharper details and cleaner frames, so motion becomes more moderate while overall quality keeps improving.
5.3 Interactive Video Generation
We focus on interactive video generation guided by camera trajectories, where the model takes a camera trajectory, an initial image, and a text prompt as inputs to generate a video that follows the defined camera motion. We fine-tune the Wan2.1-Fun-V1.1-1.3B-Control-Camera wan2025wan model with camera trajectories annotated in Sekai-Real and Sekai-Game.
5.3.1 Settings
The baseline model uses a rule-based approach to convert discrete camera control action inputs into camera poses, from which it computes Plücker embeddings sitzmann2021light for each frame and injects them into the model. In contrast, we directly use the per-frame camera poses annotated in Sekai as inputs for fine-tuning. We fine-tune the baseline model for two epochs on the drone-view portion of Sekai-Real and the entire Sekai-Game, respectively. For evaluation, we sample 50 test videos using the same procedure described in the previous section. In addition to the VBench metrics, we adopt two metrics from CameraCtrl hecameractrl: TransErr and RotErr, which quantitatively evaluate interaction following by measuring the translation and rotation discrepancies between the input trajectory and the trajectory extracted from the generated videos using Mega-SAM.
Table 2: Evaluation on interactive video generation trained on the drone-view portion of Sekai-Real. Lower is better for TransErr and RotErr; higher is better for the others.
[…]
5.3.2 Quantitative Results
Table 3: Evaluation on interactive video generation trained on Sekai-Game. Lower is better.
[…]
TransErr
($\downarrow$)
[…]
RotErr
($\downarrow$)
[…]
Table 2 shows the results of the baseline and the fine-tuned models trained on the drone-view portion of Sekai-Real. The fine-tuned model shows a clear improvement in camera control accuracy, achieving $\triangle$11.13 reduction in TransErr and $\triangle$7.31 reduction in RotErr. Meanwhile, other metrics also show consistent improvements. These results indicate that fine-tuning on Sekai not only improves interaction following but also enhances overall video generation quality in a balanced and comprehensive manner.
We also fine-tune the baseline model on Sekai-Game, as illustrated in Table 3. The fine-tuned model demonstrates consistent improvements in camera control, with the error rates reduced by more than 30% on average.